the state Pv=nRT correponds to 8gm O2 
a.pv=8rt
b.pv=rt
c.pv=rt/2

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 
PV = n RT
 
PV = w/m * RT
 
PV = 8/16 RT
 
PV = RT / 2
 
Hope it helps
 
Regards
Arun (askIITians forum expert)
Medical Student
5 years ago
formula mass of oxygen is not 32? (16 for single O atom as O2 molecule contain 2 atoms of oxygen so it isn't multiply by 2? 16*2 = 32.. 8/34